Supplement A: The owner of a pie store is considering adding a new line of pies, which will require leasing a new equipment for a monthly payment of $5,000. She estimates the variable cost for each unit produced to be approximately $4 and pies would retail for $12 each. a) How many pies does she have to sell each month in order to break even? Use both algebraic and graphic approaches to get your answer. b) How many pies must be sold to realize a profit of $3000
Answer #1

At breakeven, the proceedings from sales would be enough to meet the fixed cost

Thus if n is the breakeven then

(12-4) xn = 5000

n =5000/8 = 625

Graphical approach

Total revenue = 12n

Total cost = 4n+5000

b.

For realising the profit of 3000, the revenue should be 8000, which means

8n =8000

n=1000